OpenStudy (hesan):

solve cos45 + sin45 -tan45 - cot45

OpenStudy (anonymous):

Using the following values to solve:- \[\cos(45)=\frac{1}{\sqrt{2}}\]\[\sin(45)=\frac{1}{\sqrt{2}}\]\[\tan(45)=1\]\[\cot(45)=\frac{1}{\tan(45)}=\frac{1}{1}=1\]